Solve the differential equationdy
dy/dx + (x/1-x^2)y = x√y, y(0)=1

b) A wet porus substance in the open air loses its moisture at a rate proportional to the moisture content, if a sheet hung in the wind loses half its moisture during the first hour, then find the time when it has lost 95% moisture provided the weather
conditions remain the same.
1
Expert's answer
2020-04-09T14:17:59-0400

"dy\/dx + (x\/1-x^2)y = x\u221ay" ---(1)

"y(0)=1" ---(2)

Dividing (1) by "\sqrt y" we get;

"(1/\sqrt y)dy/dx + (x/(1-x^2))\sqrt y = x"

which is a standard Bernoulli's Differential equation which solved in the following standard manner:

Substitute "\sqrt y=t \implies 1/(2\sqrt y)*dy/dx=dt/dx" we get;

"dt\/dx+ tx\/(2(1-x^2)) = x\/2"

which is of the form : "dt\/dx+tP(x)=Q(x)"

whose solution is given as : "t(IF)=\int Q(IF)dx" where "IF=e^{\int Pdx}"

Here, "P(x)=x\/(2(1-x^2)); Q(x) = x\/2"

Thus, "IF=e^{\int xdx/(2(1-x^2))}=e^{\int 2xdx/(4(1-x^2))}"

"=e^{-(ln(1-x^2))\/4}=(1-x^2)^{-1\/4}"

Thus the solution is :

"t(1-x^2)^{-1/4}=\int xdx/(2(1-x^2)^{1/4})=\int 2xdx/(4(1-x^2)^{1/4})"

Put "1-x^2=u \implies -2xdx=du" ;we get;

"t(1-x^2)^{-1/4}=-\int du/(4u^{1/4})=(-1/3)u^{3/4}+c"

Putting t and u back; we get;

"\sqrt y(1-x^2)^{-1/4}=-(1-x^2)^{3/4}/3+c"

Using (2) we get;

"1/(1-0)=-1/3+c \implies c=4/3"

"\implies \sqrt y(1-x^2)^{-1/4}=-(1-x^2)^{3/4}/3+4/3"


b) given : "-dm\/dt=km" where k is a constant of proportionality and m is moisture content at time t

Solving we get; "m=m_oe^{-kt}; m_0" = moisture content at t=0

Also given : "m(1 hr.)=m_0e^{-k}=m_0/2\implies k=ln2"

To find: t for which "m(t)=0.05 m_0=m_0e^{-kt} \implies kt=ln20"

Thus, "t=ln20\/ln2=0.301 hr."